Hi Free Tutors:
I am having trouble with the following problme:
-2[3a-2(4a-1)]+2=2(a+1) on this problem I have tried the following steps
-6a-8a+2+2=2a+2
-14a+4=2a+2
-16a=-2
a=1/8
But according to the answer in the back of the book, answer is wrong and I am not sure what I am doing wrong.

Answer by gina(6)

-2[3a-2(4a-1)]+ 2 = 2(a+1)
-2[3a-8a+2]+2 = 2a + 1
3a - 8a + 2 = 2a + 1
-5a + 2 = 2a + 1
-5a + 5a + 2 - 1 = 2a + 5a + 1 - 1
1 = 7a
1/7 = 7a/7
1/7 = a